OpenStudy (anonymous):

Find the area of a kite with diagonals measuring 6 cm and 18 cm.

OpenStudy (anonymous):

area = (1/2)(d1*d2)= (1/2)(6*18)=54cm^2

OpenStudy (anonymous):

thanks!